Given the equation of the wave:

y(x,t)= 1.0mCos/2πm^-1 x - π/3s t)

What is

a) wavelength

b) frequency

c) angular frequency

d) period and

e) velocity


1
Expert's answer
2021-01-30T06:25:48-0500

a)

"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{2\pi\ m^{-1}}=1.0\ m."

b)

"f=\dfrac{\omega}{2\pi}=\dfrac{\dfrac{\pi}{3} \dfrac{rad}{s}}{2\pi}=0.166\ Hz."

c)

"\omega=\dfrac{\pi}{3}\ \dfrac{rad}{s}."

d)

"T=\dfrac{1}{f}=\dfrac{1}{0.166\ Hz}=6.0\ s."

e)

"v=f\lambda=0.166\ Hz\cdot1.0\ m=0.166\ \dfrac{m}{s}."